OpenStudy (anonymous):

How many units apart are the points (8, 11) and (-1, -1)?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

sqrt(8-(-1)^2+11-(-1)^2) = sqrt(9^2+12^2) = sqrt(225) = 15

OpenStudy (anonymous):

ohhh ok I see thank u !

OpenStudy (anonymous):

use distance formula